# Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting

### v2

@@ Abstract
To understand the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across four state-of-the-art large-scale MoE models released in 2025 (200B-1000B) using over 24,000 requests spanning diverse workloads. We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse future serving systems. With our insights, we then demonstrate how to improve wafer-scale GPUs as a case study, and show that minor architectural modifications leveraging the insights achieve substantial performance gains, delivering 5.3× and 3.1× average speedups on DeepSeek V3 and Qwen3 respectively. Our work presents the first comprehensive data-centric analysis of large-scale MoE models and a concrete design study using the learned lessons, with profiling traces and simulation framework already open-sourced with 1k downloads. Our traces and results are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
[…]
@@ I Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [llm_programming1, llm_programming2], translation [gpt_llm_trans1, lu2024llm_trans2], and chatbots [llm_chatbot1, llm_chatbot2]. Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [chiang2024chatbot] and the most widely used open weight models.
Unlike dense LLMs that activate all model weights uniformly, MoE models dynamically route each token to only a subset of experts, introducing substantial data movement overhead. Such overhead already exceeds 50% of execution time for small models (e.g., Mixtral 8x7B) on modest systems (2-4 GPUs), and it exacerbates further with larger models, such as DeepSeek V3 with 32 experts and 15 parameters deployed on multi-node systems (32+ GPUs) [go2025moetuner, zhang2025comet]. However, as shown in Figure 1, no prior work has systematically investigated the underlying data movement patterns in large-scale MoE models. Prior studies conducted before the emergence of large-scale MoEs [fang2025klotski, tairin2025emoe, yao2024exploiting] have confined themselves to profiling much smaller models on limited hardware, typically examining only one or two LLMs while reporting simple observations without system-level insights. As parameter size and expert count surge, new patterns have emerged but remain unexplored, leaving significant optimization opportunities on the table. Thus, a comprehensive study of the underlying data movement patterns in SOTA MoE models presents a fruitful opportunity for better efficiency.
@@
Fortunately, as we later show in the paper, MoE expert selections indeed have predictability that designers can exploit to reduce data movement. To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of four state-of-the-art MoE models ranging from 235B to 1000B parameters released in 2025. As highlighted in Figure 1, we profile DeepSeek V3 [liu2024deepseek], Llama4 Maverick [llama4], Qwen3-235B [yang2025qwen3], and Kimi K2 [team2025kimi] across 24,000 requests involving varied tasks, topics, and languages, which consumes 2000 GPU hours in total. We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files. From these extensive traces, we conduct a comprehensive analysis to uncover data movement patterns from both temporal and spatial perspectives, making our findings system-agnostic and applicable to various serving architectures at any scale. We then distill six key insights that serve as a solid foundation to understand MoE data movement and directly inform future MoE LLM serving system design, addressing critical questions that have remained unanswered in the field, such as: Is there a correlation between previously selected experts and those selected later? Are there discernible rules underlying the observed expert selection skewness? Do different tasks tend to activate different experts? Our work represents the first systematic effort to characterize data movement patterns at the scale of up-to 1000B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
To demonstrate the broad applicability of our insights, we present a case study on future GPU architecture design. Modern GPUs have adopted multi-chiplet designs due to single-die size limitations [mi300, cpelide, blackwell] and are evolving toward wafer-scale architectures enabled by emerging on-wafer packaging technologies [he2025waferlm, xu2025wsc]. By applying our insights, we develop two key innovations for future wafer-scale GPUs: (1) A task distribution algorithm with a two-level data-placement-aware command processor architecture that balances workload across dies, and (2) a data-driven predictor with hardware-managed HBM architecture that reduces inter-die communication overhead by automatically caching critical data locally. These techniques achieve an average 4.0-6.3 speedup in MoE LLM serving throughput on wafer-scale GPUs.
@@
Leveraging these insights, we conduct a case study on wafer-scale GPU architecture design. We proposed two strategies that improve MoE serving performance by 4.0-6.3 with minor hardware modifications.
@@ II-A LLM and MoE Model Architecture
Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [attention]. As shown in Figure 3(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage. During the prefill stage, all input tokens are processed simultaneously to generate the first output token. The decode stage follows immediately, where tokens are generated sequentially. The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [moe-survey]. As shown in Figure 3(b), MoE-based LLMs replace the feed-forward network (FFN) layers in traditional LLMs with MoE layers. In each layer, multiple experts are deployed, and each request is routed to a small subset of the most suitable experts based on a gating mechanism. This innovation enables MoE models to scale model parameters without incurring extra inference overhead, since only a fraction of parameters are activated for each request. However, this mechanism also introduces dynamic randomness, since expert selection is unknown until gating is completed, posing new challenges for serving systems.
@@ II-B Prior MoE Serving Systems
The MoE mechanism constitutes the primary source of data movement overhead in modern serving systems. As illustrated in Figure 2, MoE-related data movement (MoE All-to-All and MoE Weights) dominates the overhead across different DeepSeek V3 serving configurations, accounting for 60%-90% of total latency. To address this overhead, existing research on MoE serving optimization has developed numerous system-level solutions targeting different performance and cost objectives. Edge systems like MoE-Lightning [moelightning] and CoServe [suo2025coserve] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [zhang2025comet] and MegaScale-Infer [zhu2025megascale-infer] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput. Novel hardware architectures like Duplex [yun2024duplex] explore processing-in-memory to accelerate data movement in MoE LLMs.
@@ III-B1 Layer-Level
To quantify these relationships, we analyze expert pair frequencies across multiple requests and all model layers. The cumulative percentage of the most frequently selected expert pairs is shown in Figure 4(c). The results demonstrate that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of all activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively. This reveals strong, model-dependent cross-layer correlations, with Llama4 111Llama4 places dense FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers ( and ). exhibiting the strongest effect and Deepseek the weakest.
@@ III-B3 Prefill-Decode-Level Correlation
Building on the layer-level and token-level relations, we observe notable similarities in expert selection patterns between prefill and decode stages. Comparing heatmaps across stages in Figure 6(a)(b)(c)(d), we find similar distributions of bright dots, indicating that expert pairs frequently selected during prefill are also likely selected during decode. This cross-stage consistency suggests expert selection in the decode stage can be predicted using prefill data.
To quantify this similarity, we compute Spearman’s ratio () across all model layers, comparing prefill and decode heatmaps. Spearman’s Ratio measures monotonic relationships between variables, ranging from (perfect negative correlation) to (perfect positive correlation). Generally, indicates strong correlation, indicates moderate correlation, and suggests weak correlation [spearman]. The results in Figure 6(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation. This makes it possible to predict decode-stage expert selection with prefill-stage data.
@@ III-C1 Single Expert Activation Imbalance
To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [mmlu]. Figure 7(b) shows the top 10 most popular experts for each subject. Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
We further examine task impact using Chinese version of MMLU in MMLU Pro [mmlu_pro] with identical questions but different languages. Figure 7(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts. This confirms that task characteristics, including language, significantly influence expert selection, enabling task-aware serving systems that optimize expert distribution to balance workloads and reduce data movement.
@@ IV Case Study: Future GPUs Architecture Design for MoE Models
In this section, we adopt future GPU architecture design as a use case to validate our proposed insights. We follow Insight 3 to design a task allocation algorithm and leverage the temporal relation insights (Insight 1 and Insight 2) to build a data-driven predictor. We also make slight architectural modifications to support the task allocation algorithm and the data-driven predictor.
@@ IV-A Trend of Future GPU Architecture
GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations. As Moore’s Law approaches its limits [lundstrom2003moore] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [hu2023cowos], Samsung’s X-Cube [xcube], and Intel’s EMIB [emib] enable multiple chiplets within a single package. Leading vendors have adopted such designs: AMD’s MI300 [mi300] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [blackwell], and upcoming Rubin expects four [rubin].
This trend is evolving toward wafer-scale systems [cowos2]. TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [sow-x]. As shown in Figure 9(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology. Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
The TSMC SoW technology shown in Figure 9(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [lsi], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally. Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck. Remote data access requires multiple hops across die-to-die links, resulting in high latency. Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
IV-B Challenges of serving MoE with future GPUs
@@
Simplistic Task Allocation. Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs. However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [cp1, cp2]. This oblivious task-to-SM assignment generates excessive D2D traffic and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
@@ IV-C Motivation and Insights
IV-D Architecture Design
IV-D1 Architecture Overview
@@ IV-D2 Key Data Structures
PDU Prediction Table: Each PDU stores a prediction table with two key fields per expert: the cp_en bit indicating whether the expert should be cached locally (calculated by Global CP and transferred before computation), and the is_local bit tracking whether the expert is already cached (maintained by PDU itself).
IV-D3 Workflow During Kernel Launch
When a new kernel launches ( ), the Global CP implements our task allocation algorithm to split the MoE kernel into sub-kernels for each die and executes the predictor algorithm to generate duplication guidance (cp_en field is PDU). The Global CP then sends sub-kernels and prediction information to local CPs ( ). After receiving the guidance information from Global CP, each local CP allocates tasks to its SMs ( ) and configures the prediction table in the D2D controller for local HBM management ( ). After computation, the local CP collects expert duplication statistics and sends them to the Global CP to update expert distribution information.
IV-D4 Dataflow for Remote Data Access

### v4

@@ Abstract
To understand the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across four state-of-the-art large-scale MoE models released in 2025 (200B-1000B) using over 24,000 requests spanning diverse workloads. We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse serving systems. We verify these insights on both future wafer-scale GPU architectures and existing GPU systems. On wafer-scale GPUs, lightweight architectural modifications guided by our insights yield a 6.6 average speedup across four 200B–1000B models. On existing GPU systems, our insights drive the design of a prefill-aware expert placement algorithm that achieves up to 1.25 speedup on MoE computation. Our work presents the first comprehensive data-centric analysis of large-scale MoE models together with a concrete design study applying the learned lessons. Our profiling traces are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
@@ I Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [47, 76], translation [1, 41], and chatbots [11, 71]. Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [7] and the most widely used open weight models.
Unlike dense LLMs that activate all model weights uniformly, MoE models dynamically route each token to only a subset of experts, introducing substantial data movement overhead. Such overhead already exceeds 50% of execution time for small models (e.g., Mixtral 8x7B) on modest systems (2-4 GPUs), and it exacerbates further with larger models, such as DeepSeek V3 with 32 experts and 15 parameters deployed on multi-node systems (32+ GPUs) [20, 89]. However, as shown in Figure 1, no prior work has systematically investigated the underlying data movement patterns in large-scale MoE models. Prior studies conducted before the emergence of large-scale MoEs [15, 64, 84] have confined themselves to profiling much smaller models on limited hardware, typically examining only one or two LLMs while reporting simple observations without system-level insights. As parameter size and expert count surge, new patterns have emerged but remain unexplored, leaving significant optimization opportunities on the table. Thus, a comprehensive study of the underlying data movement patterns in SOTA MoE models presents a fruitful opportunity for better efficiency.
@@
Fortunately, as we later show in the paper, MoE expert selections indeed have predictability that designers can exploit to reduce data movement. To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of four state-of-the-art MoE models ranging from 235B to 1000B parameters released in 2025. As highlighted in Figure 1, we profile DeepSeek V3 [40], Llama4 Maverick [45], Qwen3-235B [82], and Kimi K2 [69] across 24,000 requests involving varied tasks, topics, and languages, which consumes 2000 GPU hours in total. We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files. From these extensive traces, we conduct a comprehensive analysis to uncover data movement patterns from both temporal and spatial perspectives, making our findings system-agnostic and applicable to various serving architectures at any scale. We then distill six key insights that serve as a solid foundation to understand MoE data movement and directly inform future MoE LLM serving system design, addressing critical questions that have remained unanswered in the field, such as: Is there a correlation between previously selected experts and those selected later? Are there discernible rules underlying the observed expert selection skewness? Do different tasks tend to activate different experts? Our work represents the first systematic effort to characterize data movement patterns at the scale of up-to 1000B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
To demonstrate the broad applicability of our insights, we present case studies on both future and existing GPU systems. On the architecture side, we observe that modern GPUs have already adopted multi-chiplet designs due to single-die size limitations [60, 10, 48] and are evolving toward wafer-scale integration enabled by emerging on-wafer packaging technologies [23, 81]. Targeting this trend, we develop a two-level data-placement-aware command processor and a hardware-managed HBM scheme that jointly balance workload across dies and reduce inter-die communication, achieving an average 6.6 speedup in MoE serving throughput on wafer-scale GPUs. On existing multi-GPU systems, we observe that prefill-stage expert selections can effectively predict decode-stage behavior. Building on this observation, we propose prefill-aware expert placement algorithms to reduce decode workload imbalance, and achieve up to 1.25 speedup.
@@
Leveraging these insights, we conduct case studies on both future and existing GPU systems. On future wafer-scale GPUs, we improve MoE throughput by 6.6 with minor hardware modifications. On existing multi-GPU systems, we achieve up to 1.25 speedup on an 8H100.
@@ II-A LLM and MoE Model Architecture
Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [72]. As shown in Figure 3(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage. During the prefill stage, all input tokens are processed simultaneously to generate the first output token. The decode stage follows immediately, where tokens are generated sequentially. The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [44]. As shown in Figure 3(b), MoE-based LLMs replace the feed-forward network (FFN) layers in traditional LLMs with MoE layers. In each layer, multiple experts are deployed, and each request is routed to a small subset of the most suitable experts based on a gating mechanism. This innovation enables MoE models to scale model parameters without incurring extra inference overhead, since only a fraction of parameters are activated for each request. However, this mechanism also introduces dynamic randomness, since expert selection is unknown until gating is completed, posing new challenges for serving systems.
@@ II-B Prior MoE Serving Systems
The MoE mechanism constitutes the primary source of data movement overhead in modern serving systems. As illustrated in Figure 2, MoE-related data movement (MoE All-to-All and MoE Weights) dominates the overhead across different DeepSeek V3 serving configurations, accounting for 60%-90% of total latency. To address this overhead, existing research on MoE serving optimization has developed numerous system-level solutions targeting different performance and cost objectives. Edge systems like MoE-Lightning [5] and CoServe [63] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [89] and MegaScale-Infer [92] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput. Novel hardware architectures like Duplex [87] explore processing-in-memory to accelerate data movement in MoE LLMs.
@@ III-B1 Layer-Level
To quantify these relationships, we analyze expert pair frequencies across requests and layers. Figure Figure 4(c) shows that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively. This reveals strong, model-dependent cross-layer correlations, with Llama4111Llama4 places dense FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers ( and ). showing the strongest effect and Deepseek the weakest.
@@ III-B3 Prefill-Decode-Level Correlation
Building on the layer-level and token-level relations, we observe notable similarities in expert selection patterns between prefill and decode stages.
Comparing heatmaps across stages in Figure 6(a)(b)(c)(d), we find similar distributions of bright dots, indicating that expert pair heatmap during prefill and decode shares similarities.
This cross-stage consistency suggests us that the prefill-collected information can guide initial decode steps until sufficient decode data accumulates.
To quantify this similarity, we compute Spearman’s ratio () across all model layers, comparing prefill and decode heatmaps. Spearman’s Ratio measures monotonic relationships between variables, ranging from (perfect negative correlation) to (perfect positive correlation). Generally, indicates strong correlation, indicates moderate correlation, and suggests weak correlation [53]. The results in Figure 6(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation. This makes it possible to predict decode-stage expert selection with prefill-stage data.
Beyond expert-pair heatmaps, we also identify prefill-to-decode correlation at the single-expert frequency level. As shown in Figure 7(a), the frequency distributions of prefill and decode stages are substantially similar, though some discrepancies exist among low-frequency experts. To examine the most popular experts, we report the overlap rate of top experts between stages in Figure 7(b): the top-5 prefill experts cover around 60% of the top-5 decode experts, rising to 75% and 90% for top-10 and top-20, respectively. This indicates that prefill information can help predict the hottest decode experts. The cross-model Spearman correlation in Figure 7(c) confirms this relationship holds across all four models.
@@ III-C1 Single Expert Activation Imbalance
To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [25]. Figure 8(b) shows the top 10 most popular experts for each subject. Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
We further examine task impact using Chinese version of MMLU in MMLU Pro [75] with identical questions but different languages. Figure 8(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts. This confirms that task characteristics, including language, significantly influence expert selection, enabling task-aware serving systems that optimize expert distribution to balance workloads and reduce data movement.
@@ IV Case Study: Future GPUs Architecture Design for MoE Models
In this section, we adopt future GPU architecture design as a use case to validate our proposed insights. We follow Insight 3 to design a task allocation algorithm and leverage the temporal relation insights (Insight 1 and Insight 2) to build a data-driven predictor. We also make slight architectural modifications to support the proposed strategies.
@@ IV-A Trend of Future GPU Architecture
GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations. As Moore’s Law approaches its limits [42] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [28], Samsung’s X-Cube [58], and Intel’s EMIB [43] enable multiple chiplets within a single package. Leading vendors have adopted such designs: AMD’s MI300 [60] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [48], and upcoming Rubin expects four [49].
This trend is evolving toward wafer-scale systems [26]. TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [57]. As shown in Figure 10(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology. Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
The TSMC SoW technology shown in Figure 10(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [27], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally. Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck. Remote data access requires multiple hops across die-to-die links, resulting in high latency. Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
IV-B Background on Wafer-scale GPU Programming model
The programming model of future wafer-scale chip remains an open question, but two major candidates have emerged: the multi-GPU-like and the single-GPU-like programming model.
Multi-GPU-like programming model. WSC-LLM [81] and MoEntwine [68] adopt a multi-GPU-like programming model that exposes the entire wafer as a multi-GPU system. Programmers can program the wafer similarly to a conventional multi-GPU system, with the key difference being the 2D mesh topology where each die communicates directly only with its neighbors. While this approach offers fine-grained control over individual dies and enables flexible software strategies, it diverges from current industry trends. For instance, although Blackwell and Rubin both integrate two compute dies, NVIDIA exposes no toolchain for fine-grained die-level control. Multi-Instance GPU (MIG) can partition a multi-die GPU into independent GPU instances, making each die act independently, but the high-speed D2D link is disabled in this mode [34], forcing inter-die communication through the 10-100 slower NVLink or PCIe and negating the benefit of multi-die packaging. Therefore, extending this programming model to wafer-scale GPUs would require substantial architectural changes to current GPU designs, including redesigning the D2D/C2C controller workflow and distributed LLC structure, making it infeasible in the near term.
Single-GPU-like programming model. HDPAT [80], Hecton [29], and our work adopt a single-GPU-like programming model that presents the entire wafer as a unified GPU, completely abstracting the multi-die topology and data placement from software. All HBM stacks form a unified memory space, and the GPU resolves data locations directly via physical addresses, making the programming experience identical to that of a monolithic GPU. We adopt this model for two key reasons. First, it aligns with commercial multi-chiplet GPUs such as Blackwell and Rubin, ensuring practical industry relevance. Second, it eliminates the programming complexity of explicit inter-die communication management through libraries such as NCCL or NVSHMEM. However, this programming simplicity transfers the optimization burden entirely to the hardware. Given the inherently distributed architecture, local versus remote data access costs vary by up to 15, yet the abstraction prevents programmers from controlling cross-die data movement. Consequently, architecture-level optimizations become critical to achieve high hardware utilization.
IV-C Challenges of serving MoE with future GPUs
@@
Simplistic Task Allocation. Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs. However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [36, 22]. This oblivious task-to-SM assignment generates excessive D2D traffic and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
@@ IV-D Motivation and Insights
IV-E Architecture Design
IV-E1 Architecture Overview
@@ IV-E2 Key Data Structures
PDU Prediction Table: Each PDU stores a prediction table with two key fields per expert: the cp_en bit indicating whether the expert should be cached locally (calculated by Global CP and transferred to each die), and the is_local bit tracking whether the expert is already cached in local HBM.
IV-E3 Workflow During Kernel Launch
When a new kernel launches ( ), the Global CP runs our task allocation algorithm to split the MoE kernel into per-die sub-kernels and executes the predictor to generate duplication guidance (cp_en field is PDU). The Global CP then sends sub-kernels and prediction information to local CPs ( ). Each local CP allocates tasks to its SMs ( ) and configures the prediction table in the D2D controller for local HBM management ( ). After computation, local CPs collect expert duplication statistics and send them to Global CP to update expert distribution information.
IV-E4 Dataflow for Remote Data Access